THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 119 — 1 October 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: A truly EPIC image from the XMM-Newton survey of the SMC. See Haberl et al. (2012).

Editorial

Dear Colleagues,

It is my pleasure to present you the 119^{th} issue of the Magellanic Clouds Newsletter. R 136 is particularly popular at the moment, featuring in the headlines of five new works; three papers concerned supernova remnants and massive evolved stars were the subject of another few papers; other new results are presented on the structure and orbit of the Magellanic Clouds, age–metallicity relation and the occurrence of Magellanic Clouds–Milky Way look-a-likes in the universe. And that’s not all.

We recommend reading Eva Grebel’s latest review on nearby metal-poor galaxies, and why not consider spending a week on the Greek island of Rhodes discussing massive stars? (See the announcement at the back of the newsletter.)

The next issue is planned to be distributed on the 1^{st} of December 2012.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Two new Wolf–Rayet stars in the LMC

Ian D. Howarth^{1} and Nolan R. Walborn^{2}

1Department of Physics & Astronomy, UCL, UK 2STScI, USA

We report the discovery of two previously unknown WN3 stars in the Large Magellanic Cloud. Both are bright (15^{th} magnitude), isolated, and located in regions covered in earlier surveys, although both are relatively weak-lined. We suggest that there may be O(10) remaining undiscovered WNE stars in the LMC.

Accepted for publication in MNRAS Available from arXiv:1207.7031

The chemistry of extragalactic carbon stars

Paul M. Woods^{1,2}, C. Walsh^{3}, M.A. Cordiner^{4} and F. Kemper^{5}

1Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK 2Jodrell Bank Centre for Astrophysics, Alan Turing Building, School of Physics and Astronomy, The University of Manchester, Oxford Road, Manchester, M13 9PL, UK 3Astrophysics Research Centre, School of Mathematics & Physics, Queen’s University Belfast, University Road, Belfast, BT7 1NN, UK 4Astrochemistry Laboratory and The Goddard Center for Astrobiology, NASA Goddard Space Flight Center, Code 691, 8800 Greenbelt Road, Greenbelt, MD 20771, USA 5Academia Sinica Institute of Astronomy & Astrophysics, P.O. Box 23-141, Taipei 10617, Taiwan

Prompted by the ongoing interest in Spitzer Infrared Spectrometer spectra of carbon stars in the Large Magellanic Cloud, we have investigated the circumstellar chemistry of carbon stars in low-metallicity environments. Consistent with observations, our models show that acetylene is particularly abundant in the inner regions of low metallicity carbon-rich AGB stars – more abundant than carbon monoxide. As a consequence, larger hydrocarbons have higher abundances at the metallicities of the Magellanic Clouds than in stars with solar metallicity. We also find the oxygen and nitrogen chemistry is suppressed at lower metallicity, as expected. Finally, we calculate molecular line emission from carbon stars in the Large and Small Magellanic Cloud and find that several molecules should be readily detectable with the Atacama Large Millimeter Array at Full Science operations.

Accepted for publication in MNRAS Available from arXiv:1207.5519

Identification of a new relatively old star cluster in the Small Magellanic Cloud

Andrés E. Piatti^{1}

1Instituto de Astronomía y Física del Espacio (IAFE), CC67, Suc 28, 1428, CABA, Argentina

We present results on the age and metallicity estimates of the astonishingly unstudied SMC cluster ESO 51-SC09, from CCD BVI photometry obtained at the ESO NTT with the EMMI attached. ESO 51-SC09 turns out to be a relatively small cluster (FWHM = 10 ± 1 pc) located ∼ 4^{∘} northward from the galaxy center. We report for the first time a mean cluster age of (7.0 ± 1.3) Gyr and a mean cluster metallicity of [Fe/H] = −1.00 ± 0.15 dex, concluding that ESO 51-SC09 belongs to the group of the oldest SMC clusters. We found that the cluster is projected onto a dominant field stellar population older (age ∼ 10–13 Gyr) and more metal-poor ([Fe/H] = −1.3 ± 0.2 dex), so that the cluster could reach its current location because of its orbital motion.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:1207.7301

Radio detection of nebulæ around four LBV stars in the LMC

C. Agliozzo^{1,2}, G. Umana^{2}, C. Trigilio^{2}, C. Buemi^{2}, P. Leto^{2}, A. Ingallinera^{1,2}, T. Franzen^{3} and A. Noriega-Crespo^{4}

1Dipartimento di Fisica e Astronomia, Sezione Astrofisica, Università degli studi di Catania, via S. Sofia 78, 95123, Catania, Italy 2INAF Osservatorio Astrofisico di Catania, via S. Sofia 78, 95123, Catania, Italy 3CSIRO Astronomy and Space Science, P.O. Box 76, Epping, NSW 1710, Australia 4Infrared Processing and Analysis Center, California Institute of Technology, Pasadena, CA 91125, USA

The nebulæ associated to four Luminous Blue Variables (LBVs) in the Large Magellanic Cloud (LMC) have been observed at 5.5 and 9 GHz using the Australia Telescope Compact Array, and radio emission has been detected for first time in these sources, R 127, R 143, S 61 and S 119. The radio maps of the nebulæ have an angular resolution of ∼ 1.′′5 and a sensitivity of 1.5–3.0 × 10^{−2} mJy beam^{−1}, and show a very similar morphology to that observed in Hα. This similarity permit us to assume that the Hα emission is not affected by strong intrinsic extinction due to dust within the nebulæ. We estimate the masses of the ionized gas in the LBVs nebulæ and their values are consistent with those measured in Galactic LBVs.

Accepted for publication in MNRAS Available from arXiv:1207.5803

SNR 0453−68.5: An asymmetric remnant and its plerion in the Large Magellanic Cloud

Randall L McEntaffer^{1}, Thomas Brantseg^{1} and Morgan Presley^{1}

1University of Iowa, Dept. of Physics & Astronomy, Van Allen Hall, Iowa City, IA 52242, USA

We present a comprehensive study of the X-ray emission from SNR 0453−68.5 in the Large Magellanic Cloud (LMC) as seen from the Chandra X-ray Observatory. This object is in a class of composite remnants that exhibit a shell of emission surrounding a central plerion, more commonly known as a pulsar wind nebula (PWN). This is one of only five remnants in the LMC with an identified PWN. We find that the shell of emission is not ejecta dominated, but rather due to shocked ISM that has been swept up by the supernova blast wave or located in a precursor cavity wall. This is supported by the morphology of the local molecular cloud as seen with the Spitzer Space Telescope. The spectral properties are consistent with a middle-aged remnant > 17000 years old. A probable point source within the central knot is determined to be the pulsar powering the synchrotron emission of the PWN. Spectral fits show the nebula is well characterized by a power law with photon index of 2.0. This index is constant over a spatial scale of 0.4–1.2 pc, which is inconsistent with younger PWN containing remnants such as the Crab Nebula and SNR 0540−69.3. These fits also contain significant contributions from an ejecta dominated thermal plasma which we interpret as evidence of mixing during an evolved interaction of the PWN with the reverse shock of the SNR. We observe no evidence that the central pulsar contains a significant velocity transverse to the line of sight and argue that despite the asymmetric surface brightness distribution the SN explosion giving birth to this remnant may have been quite symmetric.

Accepted for publication in Astrophysical Journal Available from arXiv:1202.1772

Discovery of γ-ray emission from the extragalactic pulsar wind nebula N 157B with the High Energy Stereoscopic System

H.E.S.S. Collaboration^{1}

1see http://www.mpi-hd.mpg.de/hfm/HESS/pages/collaboration/

We present the significant detection of the first extragalactic pulsar wind nebula (PWN) detected in gamma rays, N 157B, located in the large Magellanic Cloud (LMC). Pulsars with high spin-down luminosity are found to power energised nebulæ that emit gamma rays up to energies of several tens of TeV. N 157B is associated with PSR J0537−6910, which is the pulsar with the highest known spin-down luminosity. The High Energy Stereoscopic System telescope array observed this nebula on a yearly basis from 2004 to 2009 with a dead-time corrected exposure of 46 hr. The γ-ray spectrum between 600 GeV and 12 TeV is well-described by a pure power-law with a photon index of 2.8 ± 0.2 (stat) ±0.3 (syst) and a normalisation at 1 TeV of (8.2 ± 0.8 (stat) ±2.5 (syst)) ×10^{−13} cm^{−2} s^{−1} TeV^{−1}. A leptonic multi-wavelength model shows that an energy of about 4 × 10^{49} erg is stored in electrons and positrons. The apparent efficiency, which is the ratio of the TeV γ-ray luminosity to the pulsar’s spindown luminosity, 0.08% ±0.01%, is comparable to those of PWNe found in the Milky Way. The detection of a PWN at such a large distance is possible due to the pulsar’s favourable spin-down luminosity and a bright infrared photon-field serving as an inverse-Compton-scattering target for accelerated leptons. By applying a calorimetric technique to these observations, the pulsar’s birth period is estimated to be shorter than 10 ms.

Accepted for publication in A&A Letters Available from arXiv:1208.1636

The VLT–FLAMES Tarantula Survey. VI. Evidence for rotation of the young massive cluster R 136

V. Hénault-Brunet^{1}, M. Gieles^{2}, C.J. Evans^{3,1}, H. Sana^{4}, N. Bastian^{5}, J. Maíz Apellániz^{6}, W.D. Taylor^{1}, N. Markova^{7}, E. Bressert^{8,9,10}, A. de Koter^{4} and J.Th. van Loon^{11}

1Scottish Universities Physics Alliance (SUPA), Institute for Astronomy, University of Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 2Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 3UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 4Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, Postbus 94249, 1090 GE, Amsterdam, The Netherlands 5Excellence Cluster Universe, Technische Universität München, Botlzmann-Straße 2, D-85748 Garching bei München, Germany 6Instituto de Astrofísica de Andalucía–CSIC, Glorieta de la Astronomía s/n, E-18008 Granada, Spain 7Institute of Astronomy with NAO, Bulgarian Academy of Sciences, P.O. Box 136, 4700 Smoljan, Bulgaria 8School of Physics, University of Exeter, Stocker Road, Exeter EX4 4QL, UK 9European Southern Observatory, Karl-Schwarzschild-Straße 2, D-85748 Garching bei München, Germany 10Harvard-Smithsonian CfA, 60 Garden Street, Cambridge, MA 02138, USA 11Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK

Although it has important ramifications for both the formation of star clusters and their subsequent dynamical evolution, rotation remains a largely unexplored characteristic of young star clusters (few Myr). Using multi-epoch spectroscopic data of the inner regions of 30 Doradus in the Large Magellanic Cloud (LMC) obtained as part of the VLT–FLAMES Tarantula Survey, we search for rotation of the young massive cluster R 136. From the radial velocities of 36 apparently single O-type stars within a projected radius of 10 pc from the centre of the cluster, we find evidence, at the 95% confidence level, for rotation of the cluster as a whole. We use a maximum likelihood method to fit simple rotation curves to our data and find a typical rotational velocity of ∼ 3 km s^{−1}. When compared to the low velocity dispersion of R 136, our result suggests that star clusters may form with at least ∼ 20% of the kinetic energy in rotation.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1207.7071

The VLT–FLAMES Tarantula Survey. VII. A low velocity dispersion for the young massive cluster R 136

V. Hénault-Brunet^{1}, C.J. Evans^{2,1}, H. Sana^{3}, M. Gieles^{4}, N. Bastian^{5}, J. Maíz Apellániz^{6}, N. Markova^{7}, W.D. Taylor^{1}, E. Bressert^{8,9,10}, P.A. Crowther^{11} and J.Th. van Loon^{12}

1Scottish Universities Physics Alliance (SUPA), Institute for Astronomy, University of Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 2UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 3Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam, Postbus 94249, 1090 GE, Amsterdam, The Netherlands 4Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 5Excellence Cluster Universe, Technische Universität München, Boltzmann-Straße 2, D-85748 Garching bei München, Germany 6Instituto de Astrofísica de Andalucía–CSIC, Glorieta de la Astronomía s/n, E-18008 Granada, Spain 7Institute of Astronomy with NAO, Bulgarian Academy of Sciences, P.O. Box 136, 4700 Smoljan, Bulgaria 8School of Physics, University of Exeter, Stocker Road, Exeter EX4 4QL, UK 9European Southern Observatory, Karl-Schwarzschild-Straße 2, D-85748 Garching bei München, Germany 10Harvard-Smithsonian CfA, 60 Garden Street, Cambridge, MA 02138, USA 11Dept. of Physics & Astronomy, Hounsfield Road, University of Sheffield, S3 7RH, UK 12Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK

Detailed studies of resolved young massive star clusters are necessary to determine their dynamical state and evaluate the importance of gas expulsion and early cluster evolution. In an effort to gain insight into the dynamical state of the young massive cluster R 136 and obtain the first measurement of its velocity dispersion, we analyse multi-epoch spectroscopic data of the inner regions of 30 Doradus in the Large Magellanic Cloud (LMC) obtained as part of the VLT–FLAMES Tarantula Survey. Following a quantitative assessment of the variability, we use the radial velocities of non-variable sources to place an upper limit of 6 km s^{−1} on the line-of-sight velocity dispersion of stars within a projected distance of 5 pc from the centre of the cluster. After accounting for the contributions of undetected binaries and measurement errors through Monte Carlo simulations, we conclude that the true velocity dispersion is likely between 4 and 5 km s^{−1} given a range of standard assumptions about the binary distribution. This result is consistent with what is expected if the cluster is in virial equilibrium, suggesting that gas expulsion has not altered its dynamics. We find that the velocity dispersion would be ∼ 25 km s^{−1} if binaries were not identified and rejected, confirming the importance of the multi-epoch strategy and the risk of interpreting velocity dispersion measurements of unresolved extragalactic young massive clusters.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1208.0825

The sudden appearance of CO emission in LHA 115-S 65

M.E. Oksala^{1}, M. Kraus^{1}, M.L. Arias^{2,3}, M. Borges Fernandes^{4}, L. Cidale^{2,3}, M.F. Muratore^{2,3} and M. Curé^{5}

1Astronomický ústav, Akademie věd České republiky, Ond’vrejov, Czech Republic 2Departamento de Espectroscopía Estelar, Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, La Plata, Argentina 3nstituto de Astrofísica de La Plata, CCT La Plata, CONICET–UNLP, La Plata, Argentina 4Observatório Nacional, Rio de Janeiro, Brazil 5Departmento de Física y Astronomía, Facultad de Ciencias, Universidad de Valparaíso, Valparaíso, Chile

Molecular emission has been detected in several Magellanic Cloud B[e] supergiants. In this Letter, we report on the detection of CO band head emission in the B[e] supergiant LHA 115-S 65, and present a K-band near-infrared spectrum obtained with the Spectrograph for INtegral Field Observation in the Near-Infrared (SINFONI; R = 4500) on the ESO VLT UT4 telescope. The observed molecular band head emission in S 65 is quite surprising in light of a previous non-detection by McGregor, Hyland & McGinn, as well as a high resolution (R = 50 000) Gemini/Phoenix spectrum of this star taken nine months earlier showing no emission. Based on analysis of the optical spectrum by Kraus, Borges Fernandes & de Araújo, we suspect that the sudden appearance of molecular emission could be due to density build up in an outflowing viscous disk, as seen for Be stars. This new discovery, combined with variability in two other similar evolved massive stars, indicates an evolutionary link between B[e] supergiants and LBVs.

Accepted for publication in MNRAS Letters Available from arXiv:1207.6294

The age–metallicity relationship of the Large Magellanic Cloud field star population from wide-field Washington photometry

Andrés E. Piatti^{1} and Doug Geisler^{2}

1Instituto de Astronomía y Física del Espacio (IAFE), CC67, Suc 28, 1428, CABA, Argentina 2Departamente de Astronomía, Casilla 160-C, Universidad de Concapción, Concepción, Chile

We analyze age and metallicity estimates for an unprecedented database of some 5.5 million stars distributed throughout the Large Magellanic Cloud (LMC) main body, obtained from CCD Washington CT1 photometry, reported on in Piatti et al. (2012). We produce a comprehensive field star Age–Metallicity Relationship (AMR) from the earliest epoch until ∼ 1 Gyr ago. This AMR reveals that the LMC has not evolved chemically as either a closed-box or bursting system, exclusively, but as a combination of both scenarios that have varied in relative strength over the lifetime of the galaxy, although the bursting model falls closer to the data in general. Furthermore, while old and metal-poor field stars have been preferentially formed in the outer disk, younger and more metal-rich stars have mostly been formed in the inner disk, confirming an outside-in formation. We provide evidence for the formation of stars between 5 and 12 Gyr, during the cluster age gap, although chemical enrichment during this period was minimal. We find no significant metallicity gradient in the LMC. We also find that the range in the metallicity of an LMC field has varied during the lifetime of the LMC. In particular, we find only a small range of the metal abundance in the outer disk fields, whereas an average range of Δ [Fe/H] = +0.3 ± 0.1 dex appears in the inner disk fields. Finally, the cluster and field AMRs show a satisfactory match only for the last 3 Gyr, while for the oldest ages (> 11 Gyr) the cluster AMR is a remarkable lower envelope to the field AMR. Such a difference may be due to the very rapid early chemical evolution and lack of observed field stars in this regime, whereas the globular clusters are easily studied. This large difference is not easy to explain as coming from stripped ancient Small Magellanic Cloud (SMC) clusters, although the field SMC AMR is on average ∼ 0.4 dex more metal-poor at all ages than that of the LMC but otherwise very similar.

Accepted for publication in The Astronomical Journal Available from arXiv:1208.3899

Carnegie Hubble program: A mid-infrared calibration of the Hubble constant

Wendy L. Freedman^{1}, Barry F. Madore^{1}, Victoria Scowcroft^{1}, Chris Burns^{1}, Andy Monson^{1}, S. Eric Persson^{1}, Mark Seibert^{1} and Jane Rigby^{2}

1Observatories of the Carnegie Institution of Washington, USA 2Observational Cosmology Lab, NASA Goddard Space Flight Center, USA

Using a mid-infrared calibration of the Cepheid distance scale based on recent observations at 3.6 μm with the Spitzer Space Telescope, we have obtained a new, high-accuracy calibration of the Hubble constant. We have established the mid-IR zero point of the Leavitt law (the Cepheid Period–Luminosity relation) using time-averaged 3.6 μm data for ten high-metallicity, Milky Way Cepheids having independently-measured trigonometric parallaxes. We have adopted the slope of the PL relation using time-averaged 3.6 μm data for 80 long-period Large Magellanic Cloud (LMC) Cepheids falling in the period range 0.8 < log P < 1.8. We find a new reddening-corrected distance to the LMC of 18.477±0.033 (systematic) mag. We re-examine the systematic uncertainties in H_{0}, also taking into account new data over the past decade. In combination with the new Spitzer calibration, the systematic uncertainty in H_{0} over that obtained by the Hubble Space Telescope (HST) Key Project has decreased by over a factor of three. Applying the Spitzer calibration to the Key Project sample, we find a value of H_{0} = 74.3 with a systematic uncertainty of ±2.1 (systematic) km s^{−1} Mpc^{−1}, corresponding to a 2.8% systematic uncertainty in the Hubble constant. This result, in combination with WMAP7 measurements of the cosmic microwave background anisotropies and assuming a flat universe, yields a value of the equation of state for dark energy, w_{0} = −1.09 ± 0.10. Alternatively, relaxing the constraints on flatness and the numbers of relativistic species, and combining our results with those of WMAP7, Type Ia supernovæ and baryon acoustic oscillations yields w_{0} = −1.08 ± 0.10 and a value of N_{eff} = 4.13 ± 0.67, mildly consistent with the existence of a fourth neutrino species.

Accepted for publication in ApJ Available from arXiv:1208.3281

Galaxy And Mass Assembly (GAMA): In search of Milky Way–Magellanic Cloud analogues

A. Robotham^{1,2}, I.K. Baldry^{3}, J. Bland-Hawthorn^{4}, S.P. Driver^{1,2}, J. Loveday^{5}, P. Norberg^{6}, A.E. Bauer^{7}, K. Bekki^{1}, S. Brough^{7}, M. Brown^{8}, A.W. Graham^{9}, A.M. Hopkins^{7}, S. Phillipps^{10}, C. Power^{1}, A. Sansom^{11} and L.

Staveley-Smith^{1}

1International Centre for Radio Astronomy Research (ICRAR), The University of Western Australia, Australia 2University of St. Andrews, Scotland 3Liverpool John Moores University, England 4University of Sydney, Australia 5University of Sussex, England 6Durham University, England 7Australian Astronomical Observatory, Australia 8Monash University, Australia 9Swinburne University of Technology, Australia 10University of Bristol, England 11University of Central Lancashire, England

Analysing all Galaxy and Mass Assembly (GAMA) galaxies within a factor two (±0.3 dex) of the stellar mass of the Milky Way (MW), there is a 11.9% chance that one of these galaxies will have a close companion (within a projected separation of 70 kpc and radial separation of 400 km s^{−1}) that is at least as massive as the Large Magellanic Cloud (LMC). Two close companions at least as massive as the Small Magellanic Cloud (SMC) are rare at the 3.4% level. Two full analogues to the MW–LMC–SMC system were found in GAMA (all galaxies late-type and star forming), suggesting such a combination of close together, late-type, star-forming galaxies is rare: only 0.4% of MW mass galaxies (in the range where we could observe both the LMC and SMC) have such a system. In summary, the MW–LMC–SMC system is a 2.7σ event (when recast into Gaussian statistics). Using cross-correlation comparisons we find that there is a preference for SMC–LMC binary pair analogues to be located within 2 Mpc of a range of different luminosity groups. There is a particular preference for such binaries to be located near Local Group luminosity systems. When these groups are subdivided into small magnitude gap and large magnitude gap subsets, the binaries prefer to be spatially associated with the small magnitude gap systems. These systems will be dynamically less evolved, but still offer the same amount of gravitational dark matter. This suggests that binaries such as the SMC–LMC might be transient systems, usually destroyed during vigorous merger events. Details of a particularly striking analogue to the MW–SMC–LMC and M 31 complex are included.

Published in MNRAS Available from arXiv:1208.4293 and from http://adsabs.harvard.edu/abs/2012MNRAS.424.1448R

On the metallicity dependence of crystalline silicates in oxygen-rich asymptotic giant branch stars and red supergiants

O.C. Jones^{1}, F. Kemper^{2}, B.A. Sargent^{3,4}, I. McDonald^{1}, C. Gielen^{5}, Paul M. Woods^{6}, G.C. Sloan^{7}, M.L. Boyer^{4}, A.A. Zijlstra^{1}, G.C. Clayton^{8}, K.E. Kraemer^{9}, S. Srinivasan^{10} and P.M.E. Ruffle^{1}

1Jodrell Bank Centre for Astrophysics, Alan Turing Building, Oxford Road, Manchester, M13 9PL, UK 2Academia Sinica Institute of Astronomy and Astrophysics, P.O. Box 23-141, Taipei 10617, Taiwan 3Center for Imaging Science, Rochester Institute of Technology, 54 Lomb Memorial Drive, Rochester, NY 14623, USA 4Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 5Instituut voor Sterrenkunde, Katholieke Universiteit Leuven, Celestijnenlaan 200D, 3001 Leuven, Belgium 6Department of Physics and Astronomy, University College London, Gower Street, London, WC1E 6BT, UK 7Department of Astronomy, Cornell University, Ithaca, NY 14853, USA 8Department of Physics and Astronomy, Louisiana State University, Baton Rouge, LA 70803, USA 9Institute for Scientific Research, Boston College, Kenny Cottle L106B, Newton, MA 02459-1161, USA 10UPMC–CNRS UMR7095, Institut d’Astrophysique de Paris, F-75014 Paris, France

We investigate the occurrence of crystalline silicates in oxygen-rich evolved stars across a range of metallicities and mass-loss rates. It has been suggested that the crystalline silicate feature strength increases with increasing mass-loss rate, implying a correlation between lattice structure and wind density. To test this, we analyse Spitzer IRS and Infrared Space Observatory SWS spectra of 217 oxygen-rich asymptotic giant branch and 98 red supergiants in the Milky Way, the Large and Small Magellanic Clouds and Galactic globular clusters. These encompass a range of spectral morphologies from the spectrally-rich which exhibit a wealth of crystalline and amorphous silicate features to ‘naked’ (dust-free) stars. We combine spectroscopic and photometric observations with the grams grid of radiative transfer models to derive (dust) mass-loss rates and temperature. We then measure the strength of the crystalline silicate bands at 23, 28 and 33 μm. We detect crystalline silicates in stars with dust mass-loss rates which span over 3 dex, down to rates of ∼ 10^{−9} M_{⊙} yr^{−1}. Detections of crystalline silicates are more prevalent in higher mass-loss rate objects, though the highest mass-loss rate objects do not show the 23-μm feature, possibly due to the low temperature of the forsterite grains or it may indicate that the 23-μm band is going into absorption due to high column density. Furthermore, we detect a change in the crystalline silicate mineralogy with metallicity, with enstatite seen increasingly at low metallicity.

Accepted for publication in MNRAS Available from arXiv:1208.4950

The emergence of super-canonical stars in R 136-type star-burst clusters

Sambaran Banerjee^{1}, Pavel Kroupa^{1} and Seungkyung Oh^{1}

1Argelander-Institut für Astronomie, Auf dem Hügel 71, D-53121, Bonn, Germany

Among the most remarkable features of the stellar population of R 136, the central, young, massive star cluster in the 30 Doradus complex of the Large Magellanic Cloud, are the single stars whose masses substantially exceed the canonical stellar upper mass limit of 150 M_{⊙}. A recent study by us, namely, that of Banerjee, Kroupa & Oh (2012; Paper I), which involves realistic N-body computations of star clusters mimicking R 136, indicates that such ”supercanonical” (SC) stars can be formed out of a dense stellar population with a canonical initial mass function (IMF) through dynamically induced mergers of the most massive binaries. Here we study the formation of SC stars in the R 136 models of Paper I in detail. To avoid forming extraneous SC stars from initially highly eccentric primordial binaries as in Paper I, we compute additional models with only initially circular primordial binaries. We also take into account the mass evolution of the SC stars using detailed stellar evolutionary models that incorporate updated treatments of stellar winds. We find that SC stars begin to form via dynamical mergers of massive binaries from ≈ 1 Myr cluster age. We obtain SC stars with initial masses up to ≈ 250 M_{⊙} from these computations. Multiple SC stars are found to remain bound to the cluster simultaneously within a SC lifetime. These properties of the dynamically formed SC stars are consistent with those observed in R 136. In fact, the stellar evolutionary models of SC stars imply that had they formed primordially along with the rest of the R 136 cluster, i.e. violating the canonical upper limit, they would have evolved below the canonical 150 M_{⊙} limit by ≈ 3 Myr, the likely age of R 136, and would not have been observable as SC stars at the present time in R 136. This strongly supports the dynamical formation scenario of the observed SC stars in R 136.

Accepted for publication in MNRAS Available from arXiv:1208.0826

A detailed far-ultraviolet spectral atlas of O-type stars

Myron A. Smith^{1}

1Catholic University of America, USA

In this paper we present a spectral atlas covering the wavelength interval 930–1188 Å for O2–O9.5 stars using Far Ultraviolet Spectroscopic Explorer archival data. The stars selected for the atlas were drawn from three populations: Galactic main sequence (class III–V) stars, supergiants, and main sequence stars in the Magellanic Clouds, which have low metallicities. For each of these stars we have prepared FITS files comprised of paris of merged spectra for user access via the Multi-Mission Archives at Space Telescope. We chose spectra from the first population with spectral types O4, O5, O6, O7, O8, and O9.5 and used them to compile tables and figures with identifications of all possible atmospheric and ISM lines in the region 949–1188 Å. Our identified line totals for these six representative spectra are 821 (500), 992 (663), 1077 (749), 1178 (847), 1359 (1001), and 1798 (1392) lines, respectively, where the numbers in parentheses are the totals of lines formed in the atmospheres, according to spectral synthesis models. The total number of unique atmospheric identifications for the six main sequence O star template spectra is 1792, whereas the number of atmospheric lines in common to these spectra is 300. The number of identified lines decreases toward earlier types (increasing effective temperature), the whlle percentages of ”missed” features (lines not predicted from our spectral syntheses) drops from a high of 8% at type B0.2, from our recently published B star far-UV atlas, to 1–3% for type O spectra. The percentages of overpredicted lines are similar, despite their being much higher for B star spectra. We also discuss the statistics of line populations among the various elemental ionization states. Finally, as an aid to users we list those isolated lines that can be used to determine stellar temperatures and the presence of possible chemical anomalies.

Accepted for publication in Astrophysical Journal Supplements Available from http://archive.stsci.edu/prepds/fuvostars/

Magnetic field structure of the Large Magellanic Cloud from Faraday rotation measures of diffuse polarized emission

S.A. Mao^{1,2,3,4}, N.M. McClure-Griffiths^{2}, B.M. Gaensler^{5}, M. Haverkorn^{6,7}, R. Beck^{8}, D. McConnell^{2}, M. Wolleben^{9}, S. Stanimirović^{4}, J.M. Dickey^{10} and L. Staveley-Smith^{11}

1Harvard–Smithsonian Center for Astrophysics, USA 2Australia Telescope National Facility, Australia 3Jansky Fellow, National Radio Astronomy Observatory, USA 4University of Wisconsin Madison, USA 5The University of Sydney, Australia 6Radboud University, The Netherlands 7Leiden University, The Netherlands 8Max-Planck-Institut für Radioastronomie, Germany 9Square Kilometre Array South Africa, South Africa 10University of Tasmania, Australia 11International Centre for Radio Astronomy Research, Australia

We present a study of the magnetic field of the Large Magellanic Cloud (LMC), carried out using diffuse polarized synchrotron emission data at 1.4 GHz acquired at the Parkes Radio Telescope and the Australia Telescope Compact Array. The observed diffuse polarized emission is likely to originate above the LMC disk on the near side of the galaxy. Consistent negative rotation measures (RMs) derived from the diffuse emission indicate that the line-of-sight magnetic field in the LMC’s near-side halo is directed coherently away from us. In combination with RMs of extragalactic sources that lie behind the galaxy, we show that the LMC’s large scale magnetic field is likely to be of quadrupolar geometry, consistent with the prediction of dynamo theory. On smaller scales, we identify two brightly polarized filaments southeast of the LMC, associated with neutral hydrogen arms. The filaments’ magnetic field potentially aligns with the direction towards the Small Magellanic Cloud. We suggest that tidal interactions between the Small and the Large Magellanic Clouds in the past 10^{9} years is likely to have shaped the magnetic field in these filaments.

Accepted for publication in ApJ Available from arXiv:1209.1115

Detection of interstellar C_{2} and C_{3} in the Small Magellanic Cloud

Daniel E. Welty^{1}, J. Christopher Howk^{2}, Nicolas Lehner^{2} and John H. Black^{3}

1Univ. of Chicago, USA 2Univ. of Notre Dame, USA 3Chalmers Inst., Sweden

We report the detection of absorption from interstellar C_{2} and C_{3} toward the moderately reddened star Sk 143, located in the near ’wing’ region of the SMC, in optical spectra obtained with the ESO VLT/UVES. These detections of C_{2} (rotational levels J = 0–8) and C_{3} (J = 0–12) absorption in the SMC are the first beyond our Galaxy. The total abundances of C_{2} and C_{3} (relative to H_{2}) are similar to those found in diffuse Galactic molecular clouds – as previously found for CH and CN – despite the significantly lower average metallicity of the SMC. Analysis of the rotational excitation of C_{2} yields an estimated kinetic temperature T_{k} ∼ 25 K and a moderately high total hydrogen density n_{H} ∼ 870 cm^{−3} – compared to the T_{01} ∼ 45 K and n_{H} ∼ 85–300 cm^{−3} obtained from H_{2}. The populations of the lower rotational levels of C_{3} are consistent with an excitation temperature of about 34 K.

Accepted for publication in MNRAS Available from arXiv:1209.6420

Discovery of Raman-scattered lines in the massive luminous emission-line star LHA 115-S 18

Andrea F. Torres^{1,2}, Michaela Kraus^{3}, Lydia S. Cidale^{1,2}, Rodolfo Barbá^{4}, Marcelo Borges Fernandes^{5} and Estela Brandi^{1,6}

1Departamento de Espectroscopía, Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de La Plata, Paseo del Bosque S/N, La Plata, B1900FWA, Buenos Aires, Argentina 2Instituto de Astrofísica de La Plata (CCT La Plata – CONICET, UNLP), Paseo del Bosque S/N, La Plata, B1900FWA, Buenos Aires, Argentina 3Astronomický ústav, Akademie věd České republiky, Fričova 298, 251 65 Ondřejov, Czech Republic 4Instituto de Ciencias Astronómicas, de la Tierra y del Espacio (ICATE – CONICET), Casilla 467, 5400 San Juan, Argentina 5Observatório Nacional, Rua General José Cristino 77, 20921-400 São Cristovão, Rio de Janeiro, Brazil 6Comisión de Investigaciones Científicas de la Provincia de Buenos Aires (CIC), Calle 526 entre 10 y 11, Buenos Aires, Argentina

LHA 115-S 18 is a very peculiar emission-line star exhibiting the B[e] phenomenon. Located in the Small Magellanic Cloud, its spectrum shows features of an extremely wide range of excitation and ionization stages, extending from highly ionized atomic lines (Si iv, C iv, He ii) in the UV and optical regions to molecular emission bands of CO and TiO in the optical and IR regions. The most distinguishing spectral characteristic of LHA 115-S 18 is the high variability detected in the He ii λ4686 emission line, which can be a very conspicuous or completely invisible feature. In this work, we report on another peculiarity of LHA 115-S 18. From high-resolution optical spectra taken between 2000 and 2008, we discovered the appearance and strengthening of two emission features at λ6825 Å, and λ7082 Å, which we identified as Raman-scattered lines. This is the first time these lines have been detected in the spectrum of a massive luminous B[e] star. As the classification of LHA 115-S 18 is highly controversial, we discuss how the discovery of the appearance of Raman-scattered lines in this peculiar star might help us to solve this puzzle.

Accepted for publication in MNRAS Letters Available from arXiv:1209.2397

The four leading arms of the Magellanic Cloud system

M.S. Venzmer^{1,2}, J. Kerp^{1} and P.M.W. Kalberla^{1}

1Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2Institut für Astrophysik, Friedrich-Hund-Platz 1, 37077 Göttingen, Germany

The Magellanic Cloud System (MCS) interacts via tidal and drag forces with the Milky Way galaxy. Using the Parkes Galactic All-Sky Survey (GASS) of atomic hydrogen we explore the role of drag on the evolution of the so-called Leading Arm (LA). We present a new image recognition algorithm that allows us to differentiate features within a 3-D data cube (longitude, latitude, radial velocity) and to parameterize individual coherent structures. We compiled an H i object catalog of LA objects within an area of 70^{∘} × 85^{∘} (1.6 sr) of the LA region. This catalog comprises information of location, column density, line width, shape and asymmetries of the individual LA objects above the 4-σ threshold of ΔT_{b} ≃ 200 mK. We present evidence of a fourth arm segment (LA4). For all LA objects we find an inverse correlation of velocities v_{GSR} in Galactic Standard of Rest frame with Magellanic longitude. High-mass objects tend to have higher radial velocities than low-mass ones. About 1/4 of all LA objects can be characterized as head–tail (HT) structures. Using image recognition with objective criteria, it is feasible to isolate most of LA emission from the diffuse Milky Way H i gas. Some blended gas components (we estimate 5%) escape detection, but we find a total gas content of the LA that is about 50% higher than previously assumed. These methods allow the deceleration of the LA clouds to be traced towards the Milky Way disk by drag forces. The derived velocity gradient strongly supports the assumption that the whole LA originates entirely in the Large Magellanic Cloud (LMC). LA4 is observed opposite to LA1, and we propose that both arms are related, spanning about 52 kpc in space. HT structures trace drag forces even at tens of kpc altitudes above the Milky Way disk.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1208.4953

Stellar populations in the fields surrounding the LMC clusters NGC 2154 and NGC 1898

E. Chiosi^{1}, G. Baume^{2}, G. Carraro^{3,4}, E. Costa^{5} and A. Vallenari^{1}

1INAF, Osservatorio Astronomico di Padova, Vicolo Osservatorio 2, I-35122 Padova, Italy 2Facultad de Ciencias Astronómicas y Geofísicas (UNLP), Instituto de Astrofísica de la Plata (CONICET,UNLP), Paseo del Bosque s/n, La Plata, Argentina 3ESO, Alonso de Cordova 3107, Santiago de Chile, Chile 4Dipartimento di Astronomia, Universitá di Padova, Vicolo Osservatorio 3, I-35122 Padova, Italy 5Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

In this paper we present a study and comparison of the star formation rates (SFRs) in the fields around NGC 1898 and NGC 2154, two intermediate-age star clusters located in very different regions of the Large Magellanic Cloud (LMC). We also derive ages for NGC 1898, and seven minor clusters which happen to fall in the field of NGC 1898, for which basic parameters were previously unknown. We do not focus on NGC 2154, because this cluster was already investigated by Baume et al. The ages of the clusters were derived by means of the isochrone fitting method on their clean colour–magnitude diagrams. Two distinct populations of clusters were found: one cluster (NGC 2154) has a mean age of 1.7 Gyr, with indication of extended star formation over roughly a 1 Gyr period, while all the others have ages between 100 and 200 Myr. The SFRs of the adjacent fields were inferred using the downhill-simplex algorithm. Both SFRs show enhancements at 200, 400, 800 Myr, and at 1, 6 and 8 Gyr. These bursts in the SFR are probably the result of dynamical interactions between the Magellanic Clouds (MCs), and between the MCs and the Milky Way.

Accepted for publication in MNRAS Available from arXiv:1209.1311

An analysis of the FIR/Radio continuum correlation in the Small Magellanic Cloud

Howard Leverenz^{1} and Miroslav D. Filipović^{2}

1James Cook University, Townsville, QLD 7000, Australia 2University of Western Sydney, Locked Bag 1797, Penrith South DC, NSW 1797, Australia

The local correlation between far-infrared (FIR) emission and radio-continuum (RC) emission for the Small Magellanic Cloud (SMC) is investigated over scales from 3 kpc to 0.01 kpc. Here, we report good FIR/RC correlation down to ∼ 15 pc. The reciprocal slope of the FIR/RC emission correlation (RC/FIR) in the SMC is shown to be greatest in the most active star forming regions with a power law slope of ∼ 1.14 indicating that the RC emission increases faster than the FIR emission. The slope of the other regions and the SMC are much flatter and in the range of 0.63–0.85. The slopes tend to follow the thermal fractions of the regions which range from 0.5 to 0.95. The thermal fraction of the RC emission alone can provide the expected FIR/RC correlation. The results are consistent with a common source for ultraviolet (UV) photons heating dust and Cosmic Ray electrons (CRe^{−}s) diffusing away from the star forming regions. Since the CRe^{−}s appear to escape the SMC so readily, the results here may not provide support for coupling between the local gas density and the magnetic field intensity.

Published in Astrophysics & Space Science Available from arXiv:1209.0052

Discovery of new, dust-poor B[e] supergiants in the Small Magellanic Cloud

A.S. Graus^{1,2}, J.B. Lamb^{1} and M.S. Oey^{1}

1University of Michigan, USA 2University of California, Irvine, USA

We present the discovery of three new B[e] supergiants (sgB[e] stars) in the Small Magellanic Cloud (SMC). All three stars (R 15, R 38, and R 48) were identified in the course of our Runaways and Isolated O Type Star Spectroscopic Survey of the SMC (RIOTS4). The stars show optical spectra that closely resemble those of previously known B[e] stars, presenting numerous low-ionization forbidden and permitted emission lines such as [Fe ii] and Fe ii. Furthermore, our stars have luminosities of log L/L_{⊙} > 4, demonstrating that they are supergiants. However, we find lower infrared excesses and weaker forbidden emission lines than for previously identified B[e] supergiants. Thus our stars appear to either have less material in their circumstellar disks than other sgB[e] stars, or the circumstellar material has lower dust content. We suggest that these may constitute a new subclass of dust-poor sgB[e] stars.

Accepted for publication in The Astrophysical Journal Available from arXiv:1208.5486

Predicted properties of Galactic and Magellanic Classical Cepheids in the SDSS filters

Marcella Di Criscienzo^{1}, Marcella Marconi^{1}, Ilaria Musella^{1}, Michele Cignoni^{2} and Vincenzo Ripepi^{1}

1INAF–OAC, Italy 2INAF–OABo, Italy

We present the first extensive and detailed theoretical scenario for the interpretation of Cepheid properties observed in the SDSS filters. Three sets of nonlinear convective pulsation models, corresponding to the chemical compositions of Cepheids in the Milky Way, the Large Magellanic Cloud and the Small Magellanic Cloud respectively, are transformed into the SDSS bands by relying on updated model atmospheres. The resulting observables, namely the instability strip boundaries and the light curves, as well as the Period–Luminosity, the Wesenheit and the Period–Luminosity–Colour relations, are discussed as a function of the metal content, for both the fundamental and the first overtone mode. The fundamental PL relations are found to deviate from linear relations when computed over the whole observed Cepheid period range, especially at the shorter wavelenghts, confirming previous findings in the Johnson–Cousins bands. The obtained slopes are found to be mildly steeper than the ones of the semiempirical and the empirical relations available in the literature and covering roughly the same period range, with the discrepancy ranging from about 13% in u-band to about 3% in z.

Accepted for publication in MNRAS Available from arXiv:1209.4090

Multi-frequency study of supernova remnants in the Large Magellanic Cloud. Confirmation of the supernova remnant status of DEM L205

P. Maggi^{1}, F. Haberl^{1}, L.M. Bozzetto^{2}, M.D. Filipović^{2}, S.D. Points^{3}, Y.-H. Chu^{4}, M. Sasaki^{5}, W. Pietsch^{1}, R.A. Gruendl^{4}, J. Dickel^{6}, R.C. Smith^{3}, R. Sturm^{1}, E.J. Crawford^{2} and A.Y. De Horta^{2}

1Max-Planck-Institut für extraterrestrische Physik, Postfach 1312, Gießenbachstr., 85741 Garching bei München, Germany 2University of Western Sydney, Locked Bag 1797, Penrith South DC, NSW 1797, Australia 3Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, Cassilla 603 La Serena, Chile 4Astronomy Department, University of Illinois, 1002 West Green Street, Urbana, IL 61801, USA 5Institut für Astronomie und Astrophysik Tübingen, Universität Tübingen, Sand 1, 72076 Tübingen, Germany 6Physics and Astronomy Department, University of New Mexico, MSC 07-4220, Albuquerque, NM 87131, USA

The Large Magellanic Cloud (LMC) is an ideal target for the study of an unbiased and complete sample of supernova remnants (SNRs). We started an X-ray survey of the LMC with XMM–Newton, which, in combination with observations at other wavelengths, will allow us to discover and study remnants that are either even fainter or more evolved (or both) than previously known. We present new X-ray and radio data of the LMC SNR candidate DEM L205, obtained by XMM–Newton and ATCA, along with archival optical and infrared observations. We use data at various wavelengths to study this object and its complex neighbourhood, in particular in the context of the star formation activity, past and present, around the source. We analyse the X-ray spectrum to derive some remnant’s properties, such as age and explosion energy. Supernova remnant features are detected at all observed wavelengths : soft and extended X-ray emission is observed, arising from a thermal plasma with a temperature kT between 0.2 keV and 0.3 keV. Optical line emission is characterised by an enhanced [S ii]-to-Hα ratio and a shell-like morphology, correlating with the X-ray emission. The source is not or only tentatively detected at near-infrared wavelengths (shorter than 10 μm), but there is a detection of arc-like emission at mid and far-infrared wavelengths (24 and 70 μm) that can be unambiguously associated with the remnant. We suggest that thermal emission from dust heated by stellar radiation and shock waves is the main contributor to the infrared emission. Finally, an extended and faint non-thermal radio emission correlates with the remnant at other wavelengths and we find a radio spectral index between −0.7 and −0.9, within the range for SNRs. The size of the remnant is ∼ 79 × 64 pc and we estimate a dynamical age of about 35 000 years. We definitely confirm DEM L205 as a new SNR. This object ranks amongst the largest remnants known in the LMC.

The numerous massive stars and the recent outburst in star formation around the source strongly suggest that a core-collapse supernova is the progenitor of this remnant.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1208.0182

Galacto-forensics of LMC’s orbital history as a probe for the dark matter potential in the outskirts of the Galaxy

Xiaojia Zhang^{1,2}, Douglas N. C. Lin^{2,3}, Andreas Burkert^{4,5,6} and Ludwig Oser^{7}

1Department of Astronomy, Peking University, Beijing 100871, China 2Kavli Institute for Astronomy and Astrophysics, Peking University, Beijing 100871, China 3Department of Astronomy and Astrophysics, University of California, Santa Cruz, CA 95064, U.S.A. 4University Observatory Munich, Scheinerstraße 1, D-81679 Munich, Germany 5Max-Planck-Institute for Extraterrestrial Physics, Giessenbachstraße 1, 85758 Garching bei München, Germany 6Max-Planck Fellow 7Max-Planck-Institute for Astrophysics, Karl-Schwarzschild-Str. 1, D-85741 Garching bei München, Germany

The 3D observed velocities of the Large and Small Magellanic Clouds (LMC and SMC) provide an opportunity to probe the Galactic potential in the outskirts of the Galactic halo. Based on a canonical NFW model of the Galactic potential, Besla et al. (2007) reconstructed LMC and SMC’s orbits and suggested that they are currently on their first perigalacticon passage about the Galaxy. Motivated by several recent revisions of the Sun’s motion around the Galactic center, we re-examine the LMC’s orbital history and show that it depends sensitively on the dark-matter’s mass distribution beyond its present Galactic distance. We utilize results of numerical simulations to consider a range of possible structural and evolutionary models for the Galactic potentials. We find that within the theoretical and observational uncertainties, it is possible for the LMC to have had multiple perigalacticon passages on the Hubble time scale, especially if the Galactic circular velocity at the location of the Sun is greater than ∼ 228 km s^{−1}. Based on these models, a more accurate determination of the LMC’s motion may be used to determine the dark matter distribution in the outskirts of the Galactic halo.

Accepted for publication in ApJ Available from arXiv:1209.3856

On the true shape of the upper end of the stellar initial mass function: The case of R 136

Sambaran Banerjee^{1} and Pavel Kroupa^{1}

1Argelander-Institut für Astronomie, Auf dem Hügel 71, D-53121, Bonn, Germany

The shape of the stellar initial mass function (IMF) of a star cluster near its upper mass limit is a focal topic of investigation as it determines the high mass stellar content and hence the dynamics of the cluster at its embedded phase as well as during its young gas-free phase. The massive stellar content of a young cluster, however, can be substantially modified due to the dynamical ejections of the massive stars so that the present-day high-mass stellar mass function (hereafter MF) can be different than that with which the cluster is born. In the present study, we provide a preliminary estimate of this evolution of the high-mass IMF of a young cluster due to early ejections of massive stars, using the Large Magellanic Cloud massive, young cluster R 136 as an example. To that end, we utilize the results of the state-of-the-art calculations by Banerjee, Kroupa & Oh (2012) comprising direct N-body computations of realistic, binary-rich, mass-segregated models of R 136. In particular, these calculations provide the ejection fraction of stars as a function of stellar mass. We find that if the measured IMF of R 136 is granted to be canonical, as observations indicate, then the ”true” high-mass IMF of R 136 at its birth must be at least moderately top-heavy when corrected for the dynamical escape of massive stars. We conclude that the top-heaviness of the true high-mass IMF over the observationally determined one is a general feature of massive, young clusters where the dynamical ejection of massive stars is efficient. We discuss its implications and possible improvements over our current estimate.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1209.4079

The central density of R 136 in 30 Doradus

Fernando J. Selman^{1} and Jorge Melnick^{1}

1European Southern Observatory, Chile

The central density ρ_{0} of a stellar cluster is an important physical parameter to determine its evolutionary and dynamical state. The degree of mass segregation, or whether the cluster has undergone core collapse both depends on ρ_{0}. We reanalyze the results of a previous paper that gives the mass density profile of R 136 and combine them with both a conservative upper limit for the core parameter and a more uncertain recent measurement. We thus place a lower limit on ρ_{0} under reasonable and defensible assumptions about the IMF and its extrapolation to lower masses finding ρ_{0}> ∼1.5 × 10^{4} M_{⊙} pc^{−3} for the conservative assumption a < 0.4 pc for the cluster core parameter. If we use the smaller, but more uncertain value a = 0.025 pc, the central density estimate becomes larger than 10^{7} M_{⊙} pc^{−3}. A mechanism based on the destruction of a large fraction of circumstellar disks is posited to explain the hitherto unexplained increase in reddening presented in that same previous work.

Submitted to Astronomy and Astrophysics Available from arXiv:1209.3825

Review Paper

Metal-poor galaxies in the local Universe

Eva K. Grebel^{1}

1Astronomisches Rechen-Institut, Zentrum für Astronomie der Universität Heidelberg, Mönchhofstr. 12–14, 69120 Heidelberg, Germany

A galaxy’s mean metallicity is usually closely correlated with its luminosity and mass. Consequently the most metalpoor galaxies in the local universe are dwarf galaxies. Blue compact dwarfs and tidal dwarfs tend to deviate from the metallicity–luminosity relation by being too metal-poor or too metal-rich for their luminosity, respectively. A less pronounced offset separates dwarf spheroidal (dSph) and dwarf irregular galaxies, making the former too metal-rich for their luminosity, which indicates different formation conditions for these two types of dwarfs. While environment (photo-evaporation through local re-ionization by massive galaxies, tidal and ram pressure stripping) govern the observed morphology-distance relation, intrinsic properties (in particular total mass) play a decisive role in dwarf galaxy evolution with respect to the time and duration of star formation and the amount of enrichment. The metallicity distribution functions of nearby dwarfs can be understood taking pre-enrichment, gas infall, and winds into account. Many dwarfs show evidence for inhomogeneous, localized enrichment. Ultra-faint dSphs, which may have formed their metal-poor stars at high redshift via H_{2} cooling, show an overabundance of metal-deficient stars as compared to the (inner) Galactic halo, but may, along with classical dSphs, have contributed significantly to the build-up of the outer halo. The abundance ratios measured in the irregular Large Magellanic Cloud are consistent with the postulated early accretion of irregulars to form the inner Galactic halo.

Published in ”First Stars IV”, Kyoto, AIP conference proceedings, eds. K. Okamura et al. Available from arXiv:1208.2894

Announcement

”Massive Stars: From α to Ω” June 10–14, 2013, Rhodes, Greece

The Institute of Astronomy, Astrophysics, Space Applications and Remote Sensing of the National Observatory of Athens announces the conference:

Massive Stars: From α to Ω June 10–14, 2013 Rhodes, Greece

Background and Scientific Rationale

The ’Massive Stars’ meetings have enjoyed more than 40 years of startling success since the first meeting in Argentina in 1971. Held every 4 to 5 years, these meetings aim to encapsulate the current state-of-the-art of our understanding of the physics of Massive Stars and their role in the Universe. For this 10^{th} meeting in the Massive Stars series the Institute of Astronomy, Astrophysics, Space Applications and Remote Sensing of the National Observatory of Athens, invites you to the island of Rhodes, once home to one of the greatest astronomers of antiquity, Hipparchos, who is generally acknowledged as the founder of trigonometry, discoverer of precession and publisher of the first modern star catalog around 135 BC.

The conference will build on results from ongoing large-scale multi-wavelength surveys of massive stars which are being coupled with new theoretical advances dealing with stellar evolution and the processes which effect that evolution: mass loss, rotation, convection, magnetic fields, multiplicity and environment. It will tackle important problems from birth, through main sequence evolution and until core collapse.

There will be a strong focus on relating the major theoretical uncertainties afflicting stellar evolution through these phases to the current observational picture. The impetus for this focus is derived from the realization that our understanding of massive star evolution is severely challenged by new observations powered largely by technological advances in telescopes and instrumentation. This has enabled new ways of looking at old long-standing problems enabling large-scale high-quality surveys of resolved stellar populations. As theoretical approaches try to keep pace with this increase in information the cracks in our assumptions concerning stellar evolution have become more apparent, even glaring. Whereas before it might have been possible to understand some of the stars some of the time it is now clear that understanding stellar populations is a considerable challenge and will require substantial efforts to resolve.

This is an exciting time as observations have revealed large gaps in understanding of the formation and evolution of massive stars. The huge impact that massive stars have on their immediate environment, parent galaxies, and through the Universe, demands better understanding of massive star evolution from α to Ω.

Looking forward to seeing you in Rhodes!

Scientific Topics

• New observational & theoretical results from large-scale surveys (FLAMES, MiMeS, PanSTARRS, PTF), techniques (astrometry) and computation;

• Consequences of zero-age conditions on stellar evolution;

• Massive star environments, massive clusters, dynamical evolution, runaway stars and mergers;

• The importance of binaries for populations of massive stars;

• The upper end of the IMF and the role of mergers; • Massive-star magnetism and pulsation, evolutionary consequences;

• The role and evolution of stellar rotation across the H–R diagram;

• Mass loss across the H–R diagram and episodic mass-loss from LBVs and other transients;

• Constraints from endpoints;

• Massive stars at very low metallicity. Venue and Pre-registration The meeting will take place at the five-star Rodos Palace Luxury Convention Resort in the seaside town of Ixia, Rhodes, located about 4 km off downtown Rhodes, between June 10–14, 2013. We invite interested participants to visit the conference website for more information: http://a2omega.astro.noa.gr Registration begins on October 1^{st}, 2012 and the deadline for abstract submission is February 28, 2013. Scientific Organizing Committee D. Lennon (co-chair), A. Bonanos (co-chair), C. Evans, M. Hanson, R. Hirschi, E. Levesque, S. de Mink, N. Morrell, J. Puls, S. Smartt, N. Smith, G. Wade Local Organizing Committee A. Bonanos (co-chair), P. Boumis (co-chair), S. Akras, V. Antoniou, N. Castro, E. Koumpia, I. Leonidaki, Ch. Vamvatira-Nakou Contact Alceste Bonanos and Panos Boumis IAASARS, National Observatory of Athens Metaxa & Pavlou St., P. Penteli 15236 Greece Phone: +30 2108109177, 2108109162 Fax: +30 2108040453 a2omega@astro.noa.gr See also http://a2omega.astro.noa.gr